# A candidate proton cyclotron feature in the ultraluminous X-ray source NGC 4656 ULX-1

### 11 Mar 2026 v1

@@ A.2 Timing analysis
An upper limit on the pulsed fraction at the candidate frequency was estimated using the HENz2vspf task, following Cruz-Sanchez et al. (2025a, b), by comparing the observed statistic with simulated distributions. The simulations were performed assuming a sinusoidal signal with two harmonics (), and consisted of 10000 realizations for each pulsed-fraction value. From this procedure, we derive a 90% confidence upper limit of PF in the 0.3–10 keV energy range.
@@ Appendix C Chandra observation
Chandra observed NGC 4656 on April 1, 2025 (ObsID 28114) with a net exposure time of 24.5 ks, using ACIS-I in very faint mode. The data were reduced using the Chandra Interactive Analysis of Observation (CIAO 4.17) and CALDB 4.12.2, with standard procedures. The source spectrum, fitted in the 0.5-8 keV energy range, is featureless and well described by an absorbed power-law model (), resulting in the following parameters (1 errors): , , and a flux corrected for the absorption (in the 0.3-10 keV energy range for comparison with XMM–Newton results), of = erg cm-2 s-1. We then added a narrow Gaussian absorption feature to this continuum model and estimated a 3 upper limit of 0.31 keV on the line strength (fixing the line centroid energy at 3.3 keV and its width at 0.05 keV). The featureless appearance of the Chandra spectrum yields a constraint on the line normalization consistent with the absorption line strength measured by XMM–Newton, preventing any firm conclusion about its transient nature.

### 12 Mar 2026 v2

@@ A.2 Timing analysis
An upper limit on the pulsed fraction at the candidate frequency was estimated using the HENz2vspf task, following Cruz-Sanchez, N. et al. (2025, 2026), by comparing the observed statistic with simulated distributions. The simulations were performed assuming a sinusoidal signal with two harmonics (), and consisted of 10000 realizations for each pulsed-fraction value. From this procedure, we derive a 90% confidence upper limit of PF in the 0.3–10 keV energy range.
@@ Appendix C Chandra observation
Chandra observed NGC 4656 on April 1, 2025 (ObsID 28114) with a net exposure time of 24.5 ks, using ACIS-I in very faint mode. The data were reduced using the Chandra Interactive Analysis of Observation (CIAO 4.17) and CALDB 4.12.2, with standard procedures. The source spectrum, fit in the 0.5-8 keV energy range, is featureless and well described by an absorbed power-law model (), resulting in the following parameters (1 errors): , , and a flux corrected for the absorption (in the 0.3-10 keV energy range for comparison with XMM–Newton results), of = erg cm-2 s-1. We then added a narrow Gaussian absorption feature to this continuum model and estimated a 3 upper limit of 0.31 keV on the line strength (fixing the line centroid energy at 3.3 keV and its width at 0.05 keV). The featureless appearance of the Chandra spectrum yields a constraint on the line normalization consistent with the absorption line strength measured by XMM–Newton, preventing any firm conclusion about its transient nature.